# INSIGHT: A Survey of In-Network Systems for Intelligent, High-Efficiency AI and Topology Optimization

### v1

@@ 1 Introduction
In-network computation, also referred to as in-network processing, represents how computational tasks are executed within a network infrastructure. Instead of relying on end devices or dedicated servers, in-network computation involves performing computations directly on the network devices themselves, such as switches, routers, and Network Interface Cards (NICs). This approach enables applications to run within the network fabric, offering the potential for significantly higher throughput and lower latency compared to traditional models [43]. AI workloads demands on network infrastructure, encompassing both model training and inference, often involve the processing of massive datasets. A recent survey [43] indicated that over half of the data center experts indicate AI workloads to exert significant pressure on data center interconnect infrastructure in the future, surpassing even cloud computing. To meet these demands, a substantial portion of new data center facilities is expected to be dedicated to AI workloads, necessitating a re-evaluation of network architectures and capabilities.
@@
This paper contributes to the field of in-network computation for AI by providing a comprehensive analysis of the current state and future potential of integrating AI workloads with programmable network architectures. The survey systematically explores the evolution of network paradigms, such as Software-Defined Networking (SDN) [40] and Programmable Data Planes (PDPs) [27], detailing methodologies for mapping AI models onto resource-constrained network devices through efficient algorithm design and model compression techniques. The paper advances the understanding of distributed learning, particularly through in-network aggregation and federated learning, highlighting frameworks like Planter [45] and Quark [42] that simplify the development. By identifying key applications areas such as intelligent network monitoring, intrusion detection, traffic management, and Edge AI and proposing future research directions like runtime programmability and standardized benchmarks, this survey serves as a resource for researchers and practitioners aiming to develop intelligent, efficient, and scalable network infrastructures for next-generation AI applications.
@@ 2.1 Evolution of Network Architectures and the Rise of Programmable Data Planes
Building upon the foundation laid by SDN [40], the development of Programmable Data Planes (PDPs) [27] has great performance in network flexibility [43]. PDP technology enables the systematic reconfiguration of the low-level processing steps applied to network packets [29]. This programmability allows network operators to customize packet forwarding logic and to execute advanced network functions directly within the data plane, leading to improved resource utilization [43]. Modern network devices, such as switch Application-Specific Integrated Circuits (ASICs), NICs, and Field-Programmable Gate Arrays (FPGAs), leverage domain-specific languages like P4 to define and customize network protocols directly in the data plan. Using P4, operators can essentially instruct routers to conduct specified operations at terabit speeds for real-time decision-making and enhanced network visibility. This ability to flexibly reprogram the entire packet processing pipeline for offloading computations and applications to network devices is known as in-network computing [43].
@@ 2.2 The Convergence of Artificial Intelligence and Networking
The field of networking has adapted to Artificial Intelligence (AI) and its subset, machine learning (ML), to address a wide range of challenges [43]. From traffic classification and anomaly detection to network configuration and resource management, ML algorithms are being employed to analyze vast amounts of network data, identify patterns, and make intelligent decisions to optimize network performance and efficiency.
@@ 2.3 Benefits and Challenges of In-Network Computation for AI
Despite these benefits, the development and deployment of in-network AI present several significant challenges. Programmable network devices, while offering increasing flexibility, have limited resources compared to traditional servers and GPUs. These limitations include constraints on memory (SRAM, TCAM), a restricted number of processing stages, and limited computational capabilities [43]. Additionally, common ML data types like floating-point numbers are not natively supported in many data plane programming languages like P4, further complicating the implementation of complex AI models [36]. The basic arithmetic and logical operations natively supported by data plane programming languages might not be sufficient for certain AI algorithms, making complex calculations difficult [43]. Balancing model accuracy with the limited resources and scalability of in-network AI solutions remains an ongoing challenge. Updating ML models deployed within the network without disrupting traffic flow is also a complex task. Furthermore, processing encrypted traffic for in-network AI is challenging as network devices typically lack decryption capabilities. The absence of unified benchmarks makes it difficult to compare different in-network AI solutions and assess their performance and resource consumption [43]. Additionally, synchronization issues can arise in distributed in-network aggregation scenarios [16]. Overcoming these hurdles requires innovative approaches in algorithm design, model optimization, and resource management.
@@ 3.1 Mapping Machine Learning Models to Programmable Data Planes
A key step in enabling in-network AI is adapting trained ML models to operate within the architectural constraints of programmable network devices [19]. These devices typically impose limitations such as constrained memory, fixed pipeline stages, and minimal support for complex arithmetic operations. For decision tree-based models, two main approaches are used: the depth based method which maps each level of the tree to a stage in the processing pipeline, and the encode based method which transforms split criteria into binary encoded values to streamline and increase inference efficiency [22]. Different ML models require distinct deployment strategies. For instance, Binary Neural Networks (BNNs) [41], which utilize binary weights and activations, can be effectively deployed on switches using lightweight bitwise operations such as XNOR and PopCount. Other approaches include direct mapping, which translates model logic into pipeline operations; encode based mapping, which pre-processes features into discrete encodings; and lookup based methods, which enable fast inference using precomputed tables. Lookup methods are effective for models like Naive Bayes and Support Vector Machines (SVMs) that rely heavily on repetitive calculations [13].
@@ 3.2 Addressing Resource Constraints through Efficient Algorithm Design
Due to the resource constraints of network devices, such as limited memory, fixed instruction sets, and the absence of floating point support, ML models are expected to be lightweight and computation friendly [19]. Decision trees are well suited to this resource constraint environment, as they require only integer comparisons and minimal memory. In contrast, deeper architectures like conventional neural networks often go beyond hardware capabilities. To enable the proper functioning of more complex models, techniques such as fixed point arithmetic and function linearization are employed [37]. Additionally, model partitioning across pipeline stages allows for limited parallelization. A promising strategy is hybrid inference [44], where lightweight models make quick decisions in-network, and more sophisticated inference occurs in servers. This architecture maintains high throughput while preserving inference an efficiency quality.
@@ 3.3 The Role of Model Compression Techniques
Model compression is critical for deploying AI on resource constrained network devices. As model sizes continue to increase, model compression ensures deployability without sacrificing performance [31]. Common methods include pruning [8] (eliminating redundant weights), quantization [30] (reducing precision), and low-rank factorization [35] (decomposing weight matrices). These techniques collectively reduce memory footprint and inference time. More advanced approaches include knowledge distillation [20], training a model to mimic a larger model’s behavior, and neural architecture search (NAS) algorithms [46], which explore architectures optimized for specific hardware environments. These strategies yield compact yet effective models that operate within switch level constraints. Emerging innovations like conditional computation and sparse attention [9] dynamically activate only relevant portions of the model, further minimizing resource usage. When integrated with compression, these techniques significantly enhance the feasibility of deploying AI capabilities directly into the network data servers.
@@ 4.1 Accelerating Distributed AI Training with In-Network Aggregation
Training large scale AI models often requires distributing the workload across multiple machines to handle extensive datasets and computational overload [37]. Data parallelism [21] is a common strategy where each worker processes a subset of the training data and maintains a local model copy. Following each iteration, workers synchronize their model updates to ensure convergence, typically by grouping gradients across all nodes. This synchronization step is communication intensive and becomes a bottleneck as model sizes and the number of workers scale [37]. The repeated transmission of large parameter updates can saturate network capacity, leading to diminished scalability. In-network aggregation has emerged as a promising solution to this issue. By leveraging programmable network devices to perform gradient aggregation within the network itself, the communication burden on bottleneck issues can be significantly reduced, resulting in faster training, efficiency and throughput [37].
Figure 1: Comparison of the various In-Network Aggregations Frameworks
@@ 4.2 Collaborative In-Network Aggregation Strategies and Frameworks
Various frameworks were introduced to support collaborative in-network aggregation. A key challenge, however, lies in handling asynchronous gradient arrivals caused by network latency workers. For example, GOAT [17] employs a scheduling mechanism that partitions models into sub-components and distributes aggregation across multiple switches, thereby accommodating asynchronous behavior [14]. PANAMA [18] integrates a load balancing algorithm and congestion control to fairly allocate bandwidth among distributed training jobs, while Serene [5] uses hybrid synchronization to mitigate straggler effects through coordination between programmable switches and ML clients. SwitchML is most used in academic research and experimental data center environments to accelerate distributed machine learning by leveraging in-network aggregation on programmable switches [37], which exploits programmable switches to aggregate model updates in the dataplane. As it can be seen in Fig. 1, SwitchML performs the best in the speed up scenarios compared to the other frameworks like GOAT [17], PANAMA [18] and Serene[5]. SwitchML addresses hardware limitations, such as limited memory, absence of native floating point support, and constrained per-packet processing, by chunking model updates and applying quantization techniques to approximate floating point arithmetic. In addition, it introduces switch-side scoreboarding and end-host retransmissions to handle packet loss effectively.
@@ 4.3 Federated Learning and its Potential in In-Network AI
Federated Learning (FL) [14] is a decentralized approach to model training that allows numerous clients to collaborate without sharing their local data. In typical FL architectures, a central server coordinates training, and clients share only model updates. This preserves privacy and utilizes heterogeneous, decentralized data across devices such as smartphones or Internet of Things (IoT) sensors. Incorporating in-network computation into FL offers substantial performance enhancements. For instance, intermediate aggregation at edge nodes or programmable switches drastically reduces communication latency and server load [14]. This hierarchical aggregation model accelerates convergence and scales better in distributed environments.
Centralized FL can face server-side bottlenecks, while model updates, though anonymized, can inadvertently leak private information. Therefore, mechanisms like differential privacy [15] and secure multiparty computation [24] remain necessary [14]. Other issues such as data quality, device heterogeneity, computational constraints on edge devices, and privacy risks at intermediary nodes must be addressed to fully realize the potential of the federated in-network AI.
@@ 5.1 Overview of Existing Frameworks: Planter and Quark
Planter [45] Rapid Prototyping of In-Network ML Inference Tree-based (Decision Trees, Random Forests, XGBoost [7], Isolation Forest [25]), Statistical (SVM, Naive Bayes, K-means, PCA), Neural Networks, Clustering Intel Tofino/Tofino 2 [11], AMe Alveo FPGA, NVIDIA Bluefield-2 DPU [12], P4Pi (BMv2, T4P4S) [23], DELL IoT Gateway [39] Modular Design, One-Click Deployment, Supports Multiple Architectures and Targets, Automated Development Lifecycle High
Quark [42] Fully Offloading CNN Inference onto PDPs Convolutional Neural Networks (CNNs) Intel Tofino ASIC  [11] (Hardware Switch), BMv2 (Software Switch) Model Pruning, Quantization for Floating-Point Support, Modular CNN Design, Line-Line-Rate Inference with Low Latency Medium
Planter [45] is an open-source, modular framework designed for the rapid prototyping of in-network ML models across various platforms and pipeline architectures. The goal of Planter is to automate the overall process of creating, evaluating, and implementing in-network ML inference. Planter simplifies mapping trained models onto programmable data planes with little manual configuration by identifying general mapping methodologies for different ML algorithms. Several examples of supported ML models include tree-based models, statistical models, neural networks, clustering algorithms, and dimensionality reduction techniques. Planter is cross-platform, supporting a wide range of network hardware targets, including DELL IoT gateways [39], AMD Alveo FPGA [1], NVIDIA Bluefield-2 DPU [12], P4Pi (based on BMv2 and T4P4S) [23], and Intel Tofino and Tofino 2 [11]. New models, targets, architectures, and use cases can be easily integrated due to the specifics of the modular design.
@@ 5.2 Modular Design and Rapid Prototyping of In-Network ML Models
Rapid prototyping of in-network ML models is possible due to the modular design of frameworks such as Planter [45]. Hence, various combinations of ML models, datasets, and target hardware platforms can be quickly tested by researchers and developers. To offload ML classification tasks into a programmable data plane, Planter, for instance, provides one-click capabilities. The required ML model and dataset are specified by users in a configuration file. Next, Planter is responsible for other parts of the pipeline, including training the model, converting it into a network device-compatible format, and creating the P4 code required for the target architecture. Faster iteration and experimentation are possible because of this automated process, which significantly lowers the time and expertise needed to deploy ML models in-network. Planter provides automated support for every stage of the development lifecycle, including training, deployment, and testing [45]. Fig. 2 shows that In-Network computation can surpass the performance of the traditional server based systems in both latency and throughput parameters.
@@ 5.3 Cross-Platform Compatibility and Target Device Support
Cross-platform compatibility is essential for in-network AI solutions to be useful and widely adopted. By supporting a wide variety of network hardware platforms, frameworks such as Planter satisfy this requirement [45]. This enables users to test their in-network AI applications on various programmable devices, including software switches like BMv2, switch ASICs from vendors like Intel (Tofino series) [11], FPGAs from AMD [1] (Alveo series), and SmartNICs like NVIDIA Bluefield-2 [12]. These frameworks ensure a more open and accessible ecosystem for in-network AI research and development by supporting a variety of target architectures, ensuring that the developed solutions are not restricted to a particular hardware vendor or device type. This cross-platform capability is essential for the scalability and widespread deployment of in-network AI technologies in various networking environments.

### v2

@@ 1 Introduction
In-network computation, also known as in-network processing, represents how computational tasks are executed within a network infrastructure. Instead of relying on end devices or dedicated servers, in-network computation involves performing computations directly on the network devices, such as switches, routers, and Network Interface Cards (NICs). This approach enables applications to run within the network fabric, offering the potential for significantly higher throughput and lower latency compared to traditional models [38]. AI workloads demand on network infrastructure, encompassing both model training and inference, often involve the processing of massive datasets. A recent survey [38] indicated that over half of the data center experts indicate AI workloads to exert significant pressure on data center interconnect infrastructure in the future, surpassing even cloud computing. To meet these demands, a substantial portion of new data center facilities is expected to be dedicated to AI workloads, necessitating a re-evaluation of network architectures and capabilities.
@@
This paper contributes to the field of in-network computation for AI by providing a comprehensive analysis of the current state and future potential of integrating AI workloads with programmable network architectures. The survey explores the evolution of network paradigms, such as Software-Defined Networking (SDN) [24] and Programmable Data Planes (PDPs) [29], detailing methodologies for mapping AI models onto resource-constrained network devices through efficient algorithm design and model compression techniques. The paper advances the understanding of distributed learning, particularly through in-network aggregation and federated learning, highlighting frameworks like Planter [39] and Quark [42] that simplify the development. By identifying key application areas such as intelligent network monitoring, intrusion detection, traffic management, and Edge AI and proposing future research directions like runtime programmability and standardized benchmarks, this survey serves as a resource for researchers and practitioners aiming to develop intelligent, efficient, and scalable network infrastructures for next-generation AI applications.
@@ 2.1 Evolution of Network Architectures and the Rise of Programmable Data Planes
Building upon the foundation laid by SDN [24], the development of Programmable Data Planes (PDPs) [29] has great performance in network flexibility [38]. PDP technology enables the systematic reconfiguration of the low-level processing steps applied to network packets [18]. This programmability allows network operators to customize packet forwarding logic and to execute advanced network functions directly within the data plane, leading to improved resource utilization [38]. Modern network devices, such as switch Application-Specific Integrated Circuits (ASICs), NICs, and Field-Programmable Gate Arrays (FPGAs), leverage domain-specific languages like P4 to define and customize network protocols directly in the data plane. Using P4, operators can essentially instruct routers to conduct specified operations at terabit speeds for real-time decision-making and enhanced network visibility. This ability to flexibly reprogram the entire packet processing pipeline for offloading computations and applications to network devices is known as in-network computing.
@@ 2.2 The Convergence of Artificial Intelligence and Networking
The field of networking has adapted to Artificial Intelligence (AI) and its subset, machine learning (ML), to address a wide range of challenges. From traffic classification and anomaly detection to network configuration and resource management, ML algorithms are being employed to analyze vast amounts of network data, identify patterns, and make intelligent decisions to optimize network performance and efficiency.
@@ 2.3 Benefits and Challenges of In-Network Computation for AI
Despite these benefits, the development and deployment of in-network AI present several significant challenges. Programmable network devices, while offering increasing flexibility, have limited resources compared to traditional servers and GPUs. These limitations include constraints on memory (SRAM, TCAM), a restricted number of processing stages, and limited computational capabilities. Additionally, common ML data types like floating-point numbers are not natively supported in many data plane programming languages like P4, further complicating the implementation of complex AI models [21]. The basic arithmetic and logical operations natively supported by data plane programming languages might not be sufficient for certain AI algorithms, making complex calculations difficult [38]. Balancing model accuracy with the limited resources and scalability of in-network AI solutions remains an ongoing challenge. Updating ML models deployed within the network without disrupting traffic flow is also a complex task. Furthermore, processing encrypted traffic for in-network AI is challenging as network devices typically lack decryption capabilities. The absence of unified benchmarks makes it difficult to compare different in-network AI solutions and assess their performance and resource consumption. Additionally, synchronization issues can arise in distributed in-network aggregation scenarios [35]. Overcoming these hurdles requires innovative approaches in algorithm design, model optimization, and resource management.
@@ 3.1 Mapping Machine Learning Models to Programmable Data Planes
A key step in enabling in-network AI is adapting trained ML models to operate within the architectural constraints of programmable network devices [28]. These devices typically impose limitations such as constrained memory, fixed pipeline stages, and minimal support for complex arithmetic operations. For decision tree-based models, two main approaches are used: the depth-based method, which maps each level of the tree to a stage in the processing pipeline, and the encode-based method, which transforms split criteria into binary encoded values to streamline and increase inference efficiency [23]. Different ML models require distinct deployment strategies. For instance, Binary Neural Networks (BNNs) [37], which utilize binary weights and activations, can be effectively deployed on switches using lightweight bitwise operations such as XNOR and PopCount. Other approaches include direct mapping, which translates model logic into pipeline operations; encode-based mapping, which pre-processes features into discrete encodings; and lookup-based methods, which enable fast inference using precomputed tables. Lookup methods are effective for models like Naive Bayes and Support Vector Machines (SVMs) that rely heavily on repetitive calculations [14].
@@ 3.2 Addressing Resource Constraints through Efficient Algorithm Design
Due to the resource constraints of network devices, such as limited memory, fixed instruction sets, and the absence of floating point support, ML models are expected to be lightweight and computation-friendly [28]. Decision trees are well-suited to this resource-constrained environment, as they require only integer comparisons and minimal memory. In contrast, deeper architectures like conventional neural networks often go beyond hardware capabilities. To enable the proper functioning of more complex models, techniques such as fixed-point arithmetic and function linearization are employed [22]. Additionally, model partitioning across pipeline stages allows for limited parallelization. A promising strategy is hybrid inference [25], where lightweight models make quick decisions in-network, and more sophisticated inference occurs in servers. This architecture maintains high throughput while preserving inference and efficiency quality.
@@ 3.3 The Role of Model Compression Techniques
Model compression is critical for deploying AI on resource-constrained network devices. As model sizes continue to increase, model compression ensures deployability without sacrificing performance [12]. Common methods include pruning [33] (eliminating redundant weights), quantization [19] (reducing precision), and low-rank factorization [3] (decomposing weight matrices). These techniques collectively reduce memory footprint and inference time. More advanced approaches include knowledge distillation [16], training a model to mimic a larger model’s behavior, and neural architecture search (NAS) algorithms [8], which explore architectures optimized for specific hardware environments. These strategies yield compact yet effective models that operate within switch-level constraints. Emerging innovations like conditional computation and sparse attention [10] dynamically activate only relevant portions of the model, further minimizing resource usage. When integrated with compression, these techniques significantly enhance the feasibility of deploying AI capabilities directly into the network data servers.
@@ 4.1 Accelerating Distributed AI Training with In-Network Aggregation
Training large-scale AI models often requires distributing the workload across multiple machines to handle extensive datasets and computational overload [22]. Data parallelism [6] is a common strategy where each worker processes a subset of the training data and maintains a local model copy. Following each iteration, workers synchronize their model updates to ensure convergence, typically by grouping gradients across all nodes. This synchronization step is communication-intensive and becomes a bottleneck as model sizes and the number of workers scale [22]. The repeated transmission of large parameter updates can saturate network capacity, leading to diminished scalability. In-network aggregation has emerged as a promising solution to this issue. By leveraging programmable network devices to perform gradient aggregation within the network itself, the communication burden on bottleneck issues can be significantly reduced, resulting in faster training, efficiency, and throughput.
Figure 1: Comparison of Various In-Network Aggregations Frameworks
@@ 4.2 Collaborative In-Network Aggregation Strategies and Frameworks
Various frameworks were introduced to support collaborative in-network aggregation. A key challenge, however, lies in handling asynchronous gradient arrivals caused by network latency among workers. For example, GOAT [27] employs a scheduling mechanism that partitions models into sub-components and distributes aggregation across multiple switches, thereby accommodating asynchronous behavior [26]. PANAMA [15] integrates a load balancing algorithm and congestion control to fairly allocate bandwidth among distributed training jobs, while Serene [40] uses hybrid synchronization to mitigate straggler effects through coordination between programmable switches and ML clients. SwitchML is most used in academic research and experimental data center environments to accelerate distributed machine learning by leveraging in-network aggregation on programmable switches [22], which exploits programmable switches to aggregate model updates in the dataplane. As it can be seen in Fig. 1, SwitchML performs the best in the speed up scenarios compared to the other frameworks like GOAT [27], PANAMA [15], and Serene [40]. SwitchML addresses hardware limitations, such as limited memory, absence of native floating point support, and constrained per-packet processing, by chunking model updates and applying quantization techniques to approximate floating point arithmetic. In addition, it introduces switch-side scoreboarding and end-host retransmissions to handle packet loss effectively.
@@ 4.3 Federated Learning and its Potential in In-Network AI
Federated Learning (FL) [26] is a decentralized approach to model training that allows numerous clients to collaborate without sharing their local data. In typical FL architectures, a central server coordinates training, and clients share only model updates. This preserves privacy and utilizes heterogeneous, decentralized data across devices such as smartphones or Internet of Things (IoT) sensors. Incorporating in-network computation into FL offers substantial performance enhancements. For instance, intermediate aggregation at edge nodes or programmable switches drastically reduces communication latency and server load. This hierarchical aggregation model accelerates convergence and scales better in distributed environments.
Centralized FL can face server-side bottlenecks, while model updates, though anonymized, can inadvertently leak private information. Therefore, mechanisms like differential privacy [4] and secure multiparty computation [11] remain necessary [26]. Other issues, such as data quality, device heterogeneity, computational constraints on edge devices, and privacy risks at intermediary nodes, must be addressed to fully realize the potential of the federated in-network AI.
@@ 5.1 Overview of Existing Frameworks: Planter and Quark
Planter [39] Rapid Prototyping of In-Network ML Inference Tree-based (Decision Trees, Random Forests, XGBoost [5], Isolation Forest [1]), Statistical (SVM, Naive Bayes, K-means, PCA), Neural Networks, Clustering Intel Tofino/Tofino 2, AMe Alveo FPGA, NVIDIA Bluefield-2 DPU, P4Pi (BMv2, T4P4S) [17], DELL IoT Gateway Modular Design, One-Click Deployment, Supports Multiple Architectures and Targets, Automated Development Lifecycle High
Quark [42] Fully Offloading CNN Inference onto PDPs Convolutional Neural Networks (CNNs) Intel Tofino ASIC (Hardware Switch), BMv2 (Software Switch) Model Pruning, Quantization for Floating-Point Support, Modular CNN Design, Line-Line-Rate Inference with Low Latency Medium
Planter [39] is an open-source, modular framework designed for the rapid prototyping of in-network ML models across various platforms and pipeline architectures. The goal of Planter is to automate the overall process of creating, evaluating, and implementing in-network ML inference. Planter simplifies mapping trained models onto programmable data planes with little manual configuration by identifying general mapping methodologies for different ML algorithms. Several examples of supported ML models include tree-based models, statistical models, neural networks, clustering algorithms, and dimensionality reduction techniques. Planter is cross-platform, supporting a wide range of network hardware targets, including DELL IoT gateways, AMD Alveo FPGA, NVIDIA Bluefield-2 DPU, P4Pi (based on BMv2 and T4P4S) [17], and Intel Tofino and Tofino 2. New models, targets, architectures, and use cases can be easily integrated due to the specifics of the modular design.
@@ 5.2 Modular Design and Rapid Prototyping of In-Network ML Models
Rapid prototyping of in-network ML models is possible due to the modular design of frameworks such as Planter [39]. Hence, various combinations of ML models, datasets, and target hardware platforms can be quickly tested by researchers and developers. To offload ML classification tasks into a programmable data plane, Planter, for instance, provides one-click capabilities. The required ML model and dataset are specified by users in a configuration file. Next, Planter is responsible for other parts of the pipeline, including training the model, converting it into a network device-compatible format, and creating the P4 code required for the target architecture. Faster iteration and experimentation are possible because of this automated process, which significantly lowers the time and expertise needed to deploy ML models in-network. Planter provides automated support for every stage of the development lifecycle, including training, deployment, and testing [39]. Fig. 2 shows that in-network computation can surpass the performance of the traditional server-based systems in both latency and throughput parameters.
@@ 5.3 Cross-Platform Compatibility and Target Device Support
Cross-platform compatibility is essential for in-network AI solutions to be useful and widely adopted. By supporting a wide variety of network hardware platforms, frameworks such as Planter satisfy this requirement [39]. This enables users to test their in-network AI applications on various programmable devices, including software switches like BMv2, switch ASICs from vendors like Intel (Tofino series), FPGAs from AMD (Alveo series), and SmartNICs like NVIDIA Bluefield-2. These frameworks ensure a more open and accessible ecosystem for in-network AI research and development by supporting a variety of target architectures, ensuring that the developed solutions are not restricted to a particular hardware vendor or device type. This cross-platform capability is essential for the scalability and widespread deployment of in-network AI technologies in various networking environments.